OpenStudy (anonymous):

how do you simplify square root 175

OpenStudy (tkhunny):

Prime Factorization will ALWAYS solve these. 175 = 7*25 = 7*5*5 \(\sqrt{175} = \sqrt{7\cdot 25} = \sqrt{7}\sqrt{25} = 5\sqrt{7}\)